# Benchmarking pig detection and tracking under diverse and challenging conditions

### v1

@@ 6.3 Multi-object tracking results on PigTrack
Regarding association performance, the best training runs of MOTRv2 (AssA=88.7, IDF1=93.9) and MOTIP (AssA=87.9, IDF1=94.2) achieve a superior performance compared to the best SORT-based methods, ImprAsso (AssA=86.1, IDF1=92.1) and BoT-SORT (AssA=87.4, IDF1=93.5). This finding is in line with previous results on the DanceTrack dataset [49], which shares key characteristics with PigTrack, such as visually similar individuals, non-linear motion, and frequent occlusions. Under such conditions, learned association strategies appear to outperform the association heuristics commonly employed by SORT-based methods. Even when overfitting the hyperparameters of BoT-SORT to the test set (AssA=88.5, IDF1=94.5), it fails to consistently surpass the association performance of learned trackers, indicating inherent limitations of heuristic-based association strategies. The performance of the best overfitted model was included, as models in practice are often intended to operate in specific environments where overfitting is desirable.
@@
We also investigated whether the performance of the end-to-end trackers can be improved by adding more in-domain training data. Surprisingly, our results indicate that this cannot be consistently achieved across the range of data available to us. While the HOTA of MOTRv2 and MOTIP increases by 0.6 and 3.0, respectively, when the validation set is added to the training data, no notable improvement is observed when additionally incorporating the test set in a leave-one-out fashion. In fact, the HOTA of MOTRv2 even decreases slightly by 0.2 compared to the baseline trained on the training set only. This finding is surprising, as the test set was manually selected to include particularly challenging tracking scenarios that would be expected to provide valuable information during training. Further research is required to identify the exact reasons for this phenomenon. What can already be stated is that identifying the optimal amount and composition of training data for end-to-end trackers to perform well under diverse and challenging conditions appears to be a non-trivial task in itself.
6.4 Multi-object tracking results in unseen pen environment.
Table 7: Performance measures for different tracking methods on the dataset introduced by Yu et al. [42]. Best results are in bold, second-best underlined. The first block contains results taken from Yu et al. [42], while all other blocks contain results of the models from this work.
@@
Using YOLOv8 trained on data from [11]
BoT-SORT 66.3 NA NA 78.5 75.1 NA
FTO-SORT 75.6 NA NA 83.1 90.1 NA
@@
Recently, Yu et al. [42] presented FTO-SORT, a tracking system based on YOLOv8 [58] and BoT-SORT [83], designed to enhance generalization performance through several modifications. During object detection training, they employ custom data augmentations and a loss function aimed at improving the detection of challenging pig instances. For multi-object tracking, they incorporate a specialized Re-ID module, Farm Track-id Optimizer (FTO), into BoT-SORT. They trained their detector on data published by Riekert et al. [11] and evaluated tracking performance on a 2.5 min test video from a completely different pen setting (details on datasets in Tab. 2). The results indicate that the baseline (default YOLOv8 with BoT-SORT) can be vastly improved by adding the modifications of FTO-SORT. For example, the best-performing FTO-SORT model achieves a HOTA of 75.6, while the baseline only achieves a HOTA of 66.3 (Tab. 7, first block).
To evaluate the out-of-the-box performance of the tracking models from this work in an entirely new pen setting, we also applied them on the dataset introduced by Yu et al. [42] without any additional fine-tuning or hyperparameters adjustments. For a fair comparison to FTO-SORT, we evaluated the tracking models not only with the more powerful Co-DINO detector, but also with YOLOv8 trained on PigDetect, as described in Sec. 4.1.3. Looking at the performance of our YOLOv8-based trackers (Tab. 7, second block), we find that all except StrongSORT outperform FTO-SORT. For example, BoT-SORT improves upon FTO-SORT by 4.6 HOTA points. When using the same models with the more powerful Co-DINO detector (Tab. 7, third block), the difference becomes even more pronounced. For example, BoT-SORT surpasses FTO-SORT by 9.3 HOTA points. End-to-end models (Tab. 7, fourth block) perform worse than SORT-based methods on this new dataset. This contrasts with the results reported in Sec. 6.3 and suggests poorer generalization to unseen pen settings. Only when using the SORT-based models with a weaker detector (YOLOv8), the end-to-end methods can compete with them. Nonetheless, both end-to-end models still outperform FTO-SORT.
The results obtained on the dataset of Yu et al. [42] convey an important insight: To obtain powerful and generalizable models in practice, a lot can already be achieved by training and hyperparameter-tuning general-purpose models on carefully curated, diverse datasets. While further improvements via incremental architecture changes or other modifications are certainly possible, we argue that establishing strong baselines in a data-centric manner is a crucial first step for all further analyses and comparisons.
@@ 6.5 Multi-object tracking oracle analysis on PigTrack
For all methods, association performance on the PigTrack test set improves substantially when oracle detections are provided (Tab. 8). For instance, both StrongSORT and BoT-SORT achieve an AssA of 98.0 and an IDF1 of 98.9. Similarly strong association performance has been reported for oracle analyses on the MOT17 pedestrian dataset [57], as shown by Sun et al. [49], suggesting that detection performance is often a bottleneck for SORT-based methods. These analyses suggest that SORT-based methods still have strong potential in case that object detection can be substantially improved. Whether this potential can be fully realized in practice using image-based detectors remains questionable, especially under challenging conditions such as extreme occlusions, poor visibility or motion blur, which are prevalent in PigTrack. The results for MOTRv2 and MOTIP (see Sec. 6.3) further suggest that superior association performance can be achieved despite inferior detections. This suggests that good detections are not necessarily a prerequisite for good association performance.
@@ 6.6 Error analysis of end-to-end trainable tracking models
Figure 5: Characteristic failure cases of end-to-end trainable trackers. In both subfigures, the predicted IDs in frame t and frame t+1 are indicated by the color of the bounding boxes. If the predicted ID is the same in both frames, the bounding box has the same color. (a) shows two examples for MOTRv2. The two-colored dotted bounding boxes in frame t represent overlapping duplicate predictions made by the model in case of occlusions. The alternating colors indicate that the model produced two distinct bounding boxes with different predicted IDs at the same location. In frame t+1, the corresponding two IDs are not predicted at all. (b) shows two examples for MOTIP. The dotted bounding boxes indicate where errors happen. For instance, in the top row, the central pig is incorrectly assigned a different ID in frame t+1.
@@
MOTRv2 uses track queries to predict bounding boxes for each frame of the input video, where distinct objects are represented by distinct track queries (see Sec. 5.1.2). A common failure case emerges when one tracked objects severely occludes another one. In this case, the track query of the occluded object often latches onto the foreground object instead of not making any prediction. As a result, both track queries often produce near-identical bounding boxes (Fig. 5a, frame t). Following the duplicate predictions, we often observe on the subsequent frames that neither of the two track queries latches onto an object anymore, leading to odd-looking cases where pigs are not detected at all although they are clearly visible (Fig. 5a, frame t+1). We conjecture that this is due to the fact that duplicate predictions of the same object are penalized during training. If duplicate predictions still happen at inference time, this information can be retrieved via self-attention between track queries, and then be used to reduce the score of the affected bounding boxes. Similar errors have also been reported on the MOT17 pedestrian dataset [57] by Zhang et al. [47]. Interestingly, they found that this issue does not occur on the DanceTrack dataset [49] and attributed this to the benefits of a larger training set. MOT17 has $5316$ frames in the training set, while DanceTrack has 41 7964179641\,79641 796. In our study, increasing the training set size from $9959$ to roughly 26 0002600026\,00026 000 frames (see Sec. 6.3) did not consistently mitigate the issue.
@@
MOTIP uses the ID predictions of bounding boxes on previous frames as information to predict IDs for all bounding boxes on the current frame (see Sec. 5.1.2). Surprisingly, incorrect ID predictions often emerge in simple scenarios, where classical association heuristics employed by SORT-based models would not produce any errors. For example, even when the bounding boxes on subsequent frames overlap almost perfectly (Fig. 5b top) or the pigs move in a simple linear fashion (Fig. 5b bottom), IDs might be predicted incorrectly. One potential reason for these kinds of failures are ID conflicts due to duplicate ID predictions. To count the number of ID conflicts, we identified for each ID the number of bounding boxes that predict this ID with a minimum score of Xeven though another bounding box also predicts it with X%. From this number, we then subtract one since an ID is supposed to be predicted one time. For example, if three bounding boxes predict the ID 15 with a score of at least X%, this would be counted as two ID conflicts. By running this analysis on the entire PigTrack dataset, the proportion of ID conflicts can be calculated (Fig. 6). In total, roughly 1.8% of bounding boxes predict an ID that has already been assigned to another box with a score of at least 0.5. Even if a score of 0.95 is required for an ID conflict, roughly 1% of all bounding boxes fulfil this criterion. Given that a pen in our dataset contains between 7 and 26 pigs, this means that such an error occurs on average every couple of frames. Since IDs can only be assigned uniquely, those duplicate ID predictions with the lowest confidence will in most cases get assigned a new ID, leading to the errors that are observed in Fig. 5b. Similarly, the large amount of ID switches observed for MOTIP (Tab. 6) can be explained by this. Interestingly, these characteristic errors of MOTIP were not reported on the DanceTrack dataset by Gao et al. [48], and we were able to reproduce this finding. One possible explanation for this discrepancy is the abundance of in-domain data in the pretraining corpus: DanceTrack features videos of humans, and COCO includes a large number of images with humans. In contrast, COCO does not contain images with pigs. Additionally, DanceTrack is larger then PigTrack, potentially enabling more effective training, and has a higher frame rate, which may help reduce errors at inference time.
@@ 7 Conclusion
In this work, we carefully curated two benchmark datasets, PigDetect for object detection and PigTrack for multi-object tracking, and evaluated several recent methods for both tasks. To our knowledge, this is the first work in the context of pig farming that focuses on the systematic acquisition of diverse and challenging benchmark datasets for detection and tracking. Such datasets are crucial prerequisites for method development and comparison in the context of animal farming. Our benchmarking results indicate that the datasets are able to differentiate between the evaluated methods, often highlighting particular strengths and weaknesses, which is an essential characteristic of a benchmark dataset.
Regarding object detection, substantial performance gains be achieved by employing more optimized state-of-the-art models instead of real-time detectors, at the cost of slower inference. In the end, practitioners need to select the right model based on the desired performance-speed trade-off for the application at hand. Regarding multi-object tracking, we found that SORT-based models tend to achieve a better detection performance, while end-to-end trackers are better at association. Although better association performance is arguably more important than detection quality, end-to-end models come with additional drawbacks: Both annotation and training require large amounts of resources and the generalization to new pens is worse than for SORT-based models. For this reason, we do not currently consider them to be preferable to SORT-based methods. However, with the development of new end-to-end tracking methods in the future, the benefits might eventually outweigh their drawbacks. Given the relatively limited visual variability between different pigs and pen settings, the emergence of robust and broadly applicable end-to-end or SORT-based pig tracking models appears to be a realistic prospect. In fact, our trained detection models and the trackers based on it already achieved good performance in unseen pen settings. The benchmark datasets developed in this work can serve as a tool to constantly monitor new developments and identify models that improve upon the approaches that are currently dominating.
@@
Going forward, the detection and tracking functionality introduced in this work can be used as a modular component in more sophisticated action understanding frameworks. For example, the capabilities of vision-language models [100, 101] that can be prompted with spatial localizations, such as bounding boxes, should be continuously monitored. These models might offer a convenient way to extract action-related information from images or videos in the future. Similarly, trainable models for spatio-temporal action understanding [19, 20, 21] have not been thoroughly explored yet. For this purpose, our datasets might also be extended with spatio-temporal action labels. Since challenging tracking scenarios naturally include interesting actions such as mounting, playing or fighting, PigTrack is well suited for extending annotations to the action domain. All trained models, the source code for training and inference, as well as the PigDetect and PigTrack datasets are made publicly available. This way, researchers and practitioners can reproduce our results, use our models for downstream analyses, or add their own training data to obtain customized models for their setting. We aim to encourage open data and open source principles to enable a more holistic scientific process in the context of animal farming.

### v2

@@ 6.4 Multi-object tracking results on PigTrack
Regarding association performance, the best training runs of MOTRv2 (AssA=88.7, IDF1=93.9) and MOTIP (AssA=87.9, IDF1=94.2) achieve a superior performance compared to the best SORT-based methods, ImprAsso (AssA=86.1, IDF1=92.1) and BoT-SORT (AssA=87.4, IDF1=93.5). This finding is in line with previous results on the DanceTrack dataset [sun2022dancetrack], which shares key characteristics with PigTrack, such as visually similar individuals, non-linear motion, and frequent occlusions. Under such conditions, learned association strategies appear to outperform the association heuristics commonly employed by SORT-based methods. Even when overfitting the hyperparameters of BoT-SORT to the test set (AssA=88.5, IDF1=94.5), it fails to consistently surpass the association performance of learned trackers, indicating inherent limitations of heuristic-based association strategies. The performance of the best overfitted model was included, as models in practice are often intended to operate in specific environments where overfitting is desirable.
@@
We also investigated whether the performance of the end-to-end trackers can be improved by adding more in-domain training data. Surprisingly, our results indicate that this cannot be consistently achieved across the range of data available to us. While the HOTA of MOTRv2 and MOTIP increases by 0.6 and 3.0, respectively, when the validation set is added to the training data, no notable improvement is observed when additionally incorporating the test set in a leave-one-out fashion. In fact, the HOTA of MOTRv2 even decreases slightly by 0.2 compared to the baseline trained on the training set only. This finding is surprising, as the test set was manually selected to include particularly challenging tracking scenarios that would be expected to provide valuable information during training. Further research is required to identify the exact reasons for this phenomenon. Determining the optimal amount and composition of training data for end-to-end trackers appears to be a non-trivial task.
6.5 Multi-object tracking results on third-party dataset
Table 8: Performance measures for different tracking methods on the dataset introduced by yu2025fto. Best results are in bold, second-best underlined. The first block contains results taken from yu2025fto, while all other blocks contain results of the models from this work.
@@ 6.5 Multi-object tracking results on third-party dataset
Using YOLOv8 trained on data from [riekert_automatically_2020]
BoT-SORT 66.3 - - 78.5 75.1 -
FTO-SORT 75.6 - - 83.1 90.1 -
@@
Recently, yu2025fto presented FTO-SORT, a tracking system based on YOLOv8 [Jocher_Ultralytics_YOLO_2023] and BoT-SORT [aharon2022bot], designed to enhance generalization performance through several modifications. During object detection training, they employ custom data augmentations and a loss function aimed at improving the detection of challenging pig instances. For multi-object tracking, they incorporate a specialized Re-ID module, Farm Track-id Optimizer (FTO), into BoT-SORT. They trained their detector on data published by riekert_automatically_2020 and evaluated tracking performance on a 2.5 min test video from a completely different pen setting (details on datasets in Tab. 2). The results indicate that the baseline (default YOLOv8 with BoT-SORT) can be vastly improved by adding the modifications of FTO-SORT. For example, the best-performing FTO-SORT model achieves a HOTA of 75.6, while the baseline only achieves a HOTA of 66.3 (Tab. 8, first block).
To evaluate the out-of-the-box performance of the tracking models from this work in an entirely new pen setting, we also applied them on the dataset introduced by yu2025fto without any additional fine-tuning or hyperparameters adjustments. For a fair comparison to FTO-SORT, we evaluated the tracking models not only with the more powerful Co-DINO detector, but also with YOLOv8 trained on PigDetect, as described in Sec. 4.1.3. Looking at the performance of our YOLOv8-based trackers (Tab. 8, second block), we find that all except StrongSORT outperform FTO-SORT. For example, BoT-SORT improves upon FTO-SORT by 4.6 HOTA points. When using the same models with the more powerful Co-DINO detector (Tab. 8, third block), the difference becomes even more pronounced. For example, BoT-SORT surpasses FTO-SORT by 9.3 HOTA points. End-to-end models (Tab. 8, fourth block) perform worse than SORT-based methods on this new dataset. This contrasts with the results reported in Sec. 6.4 and suggests poorer generalization to unseen pen settings. Only when using the SORT-based models with a weaker detector (YOLOv8), the end-to-end methods can compete with them. Nonetheless, both end-to-end models still outperform FTO-SORT.
The results obtained on the dataset of yu2025fto convey an important insight: To obtain powerful and generalizable models in practice, a lot can already be achieved by training and hyperparameter-tuning general-purpose models on carefully curated, diverse datasets. While further improvements via incremental architecture changes or other modifications are certainly possible, we argue that establishing strong baselines in a data-centric manner is a crucial first step for all further analyses and comparisons.
@@ 6.6 Multi-object tracking oracle analysis on PigTrack
For all methods, association performance on the PigTrack test set improves substantially when oracle detections are provided (Tab. 9). For instance, both StrongSORT and BoT-SORT achieve an AssA of 98.0 and an IDF1 of 98.9. Similarly strong association performance has been reported for oracle analyses on the MOT17 pedestrian dataset [milan2016mot16], as shown by sun2022dancetrack, suggesting that detection performance is often a bottleneck for SORT-based methods. These analyses suggest that SORT-based methods still have strong potential in case that object detection can be substantially improved. Whether this potential can be fully realized in practice using image-based detectors remains questionable, especially under challenging conditions such as extreme occlusions, poor visibility or motion blur, which are prevalent in PigTrack. The results for MOTRv2 and MOTIP (see Sec. 6.4) further suggest that superior association performance can be achieved despite inferior detections. This suggests that good detections are not necessarily a prerequisite for good association performance.
@@ 6.7 Error analysis of end-to-end trainable tracking models
Figure 5: Characteristic failure cases of end-to-end trainable trackers. In both subfigures, the predicted IDs in frame t and frame t+1 are indicated by the color of the bounding boxes. If the predicted ID is the same in both frames, the bounding box has the same color. (a) shows two examples for MOTRv2. The two-colored dotted bounding boxes in frame t represent overlapping duplicate predictions made by the model in case of occlusions. In other words, in these cases the model produces two distinct bounding boxes with different predicted IDs at the same location. In frame t+1, the corresponding two IDs are not predicted at all. (b) shows two examples for MOTIP. The dotted bounding boxes indicate where errors happen. For instance, in the top row, the central pig is incorrectly assigned a different ID in frame t+1 even though it barely moved.
@@
MOTRv2 uses track queries to predict bounding boxes for each frame of the input video, where distinct objects are represented by distinct track queries (see Sec. 5.1.2). A common failure case emerges when one tracked objects severely occludes another one. In this case, the track query of the occluded object often latches onto the foreground object instead of not making any prediction. As a result, both track queries often produce near-identical bounding boxes (Fig. 5a, frame t). Following the duplicate predictions, we often observe on the subsequent frames that neither of the two track queries latches onto an object anymore, leading to odd-looking cases where pigs are not detected at all although they are clearly visible (Fig. 5a, frame t+1). We conjecture that this is due to the fact that duplicate predictions of the same object are penalized during training. If duplicate predictions still happen at inference time, this information can be retrieved via self-attention between track queries, and then be used to reduce the score of the affected bounding boxes. Similar errors have also been reported on the MOT17 pedestrian dataset [milan2016mot16] by zhang_motrv2_2023. Interestingly, they found that this issue does not occur on the DanceTrack dataset [sun2022dancetrack] and attributed this to the benefits of a larger training set. MOT17 has $5316$ frames in the training set, while DanceTrack has 41 79641\,796. In our study, increasing the training set size from $9959$ to roughly 26 00026\,000 frames (see Sec. 6.4) did not consistently mitigate the issue.
@@
MOTIP uses the ID predictions of bounding boxes on previous frames as information to predict IDs for all bounding boxes on the current frame (see Sec. 5.1.2). Surprisingly, incorrect ID predictions often emerge in simple scenarios, where classical association heuristics employed by SORT-based models would not produce any errors. For example, even when the bounding boxes on subsequent frames overlap almost perfectly (Fig. 5b top) or the pigs move in a simple linear fashion (Fig. 5b bottom), IDs might be predicted incorrectly. One potential reason for these kinds of failures are ID conflicts due to duplicate ID predictions. To count the number of ID conflicts, we identified for each ID the number of bounding boxes that predict this ID with a minimum score of $s_{\min}$ even though another bounding box also predicts it with $s_{\min}$. From this number, we then subtract one since an ID is supposed to be predicted one time. For example, if three bounding boxes predict the ID 15 with a score of at least $s_{\min}$, this would be counted as two ID conflicts. The minimum ID score $s_{\min}$ is a variable, which we varied from 0.5 to 0.95 in steps of 0.05. By running this analysis on the entire PigTrack dataset, the proportion of ID conflicts can be calculated for the different minimum ID scores $s_{\min}$ (Fig. 6). In total, roughly 1.8% of bounding boxes predict an ID that has already been assigned to another box with a score of at least 0.5. Even if a score of 0.95 is required for an ID conflict, roughly 1% of all bounding boxes fulfil this criterion. Given that a pen in our dataset contains between 7 and 26 pigs, this means that such an error occurs on average every couple of frames. Since IDs can only be assigned uniquely, those duplicate ID predictions with the lowest confidence will in most cases get assigned a new ID, leading to the errors that are observed in Fig. 5b. Similarly, the large amount of ID switches observed for MOTIP (Tab. 7) can be explained by this. Interestingly, these characteristic errors of MOTIP were not reported on the DanceTrack dataset by gao2025multiple, and we were able to reproduce this finding. One possible explanation for this discrepancy is the abundance of in-domain data in the pretraining corpus: DanceTrack features videos of humans, and COCO includes a large number of images with humans. In contrast, COCO does not contain images with pigs. Additionally, DanceTrack is larger then PigTrack, potentially enabling more effective training, and has a higher frame rate, which may help reduce errors at inference time.
@@ 7 Conclusion
In this work, we curated two benchmark datasets based on footage from conventional indoor barn environments: PigDetect for object detection and PigTrack for multi-object tracking. We then investigated the performance of several recent general-purpose methods on both datasets. To our knowledge, this is the first work in the context of pig farming that focuses on the systematic acquisition of diverse and challenging benchmark datasets for detection and tracking. Such datasets are crucial prerequisites for method development and comparison in animal farming. Our benchmarking results indicate that the datasets are able to differentiate between the evaluated methods, often highlighting particular strengths and weaknesses, which is an essential characteristic of a benchmark dataset.
Regarding object detection, substantial performance gains can be achieved by employing more optimized state-of-the-art models instead of real-time detectors, at the cost of slower inference. In the end, practitioners need to select the right model based on the desired performance-speed trade-off for the application at hand. Regarding multi-object tracking, we found that SORT-based models tend to achieve a better detection performance, while end-to-end trackers are better at association. Although better association performance is arguably more important than detection quality, end-to-end models come with additional drawbacks: Both annotation and training require large amounts of resources and the generalization to new pens is worse than for SORT-based models. For this reason, we do not currently consider them to be preferable to SORT-based methods. However, with the development of new end-to-end tracking methods in the future, the benefits might eventually outweigh their drawbacks. Given the relatively limited visual variability between different pigs and pen settings, the emergence of robust and broadly applicable end-to-end or SORT-based pig tracking models appears to be a realistic prospect. In fact, our trained detection models and the trackers based on it already achieved good performance in unseen pen settings. The benchmark datasets developed in this work can serve as a tool to constantly monitor new developments and identify models that improve upon the approaches that are currently dominating.
@@
While the models trained on our datasets demonstrate promising results in unseen pen settings, several limitations must be acknowledged. (1) Although our datasets contain many challenging examples, this does not imply flawless performance in such cases. Scenarios like poor visibility or heavily overlapping animals might still lead to model errors, even in pen environments similar to those used during training. (2) The datasets that were used to evaluate test performance are limited in size and restricted to conditions that are relatively similar to the training data. Specifically, the majority of images and videos in our datasets originate from conventional indoor pens with common crossbreeds. Therefore, models might perform worse for outdoor pens, pens with unusual interiors or pig breeds, or substantially different camera perspectives. Future work should focus on combining and extending existing datasets to build a more comprehensive benchmark, and enable a more robust assessment of model performance across diverse settings. (3) Examples involving heavy piling and occlusions, where unambiguous annotation was not possible, were not included in our datasets. This illustrates the practical limitations of detection and tracking and highlights the need for further research on how to address such situations when using computer vision for individualized behavioral analysis. For example, it might be possible to leverage ear tag detection methods to recover lost tracks resulting from complex scenarios that current tracking models cannot handle. Re-identification via ear tags might also be useful in situations where a single camera cannot cover the entire pen.
Going forward, the detection and tracking functionality introduced in this work can be used as a modular component in more sophisticated action understanding frameworks. For example, the capabilities of vision-language models [you2023ferret, zhang2024ferret] that can be prompted with spatial localizations, such as bounding boxes, should be continuously monitored. These models might offer a convenient way to extract action-related information from images or videos in the future. Similarly, trainable models for spatio-temporal action understanding [gritsenko_end--end_2023-1, ryali2023hiera, wu_memvit_2022] have not been thoroughly explored yet. For this purpose, our datasets might also be extended with spatio-temporal action labels. Since challenging tracking scenarios naturally include interesting actions such as mounting, playing or fighting, PigTrack is well suited for extending annotations to the action domain. All trained models, the source code for training and inference, as well as the PigDetect and PigTrack datasets are made publicly available. This way, researchers and practitioners can reproduce our results, use our models for downstream analyses, or add their own training data to obtain customized models for their setting. We aim to encourage open data and open source principles to enable a more holistic scientific process in the context of animal farming.